# Dark Matter Search Results from 4.2 Tonne-Years of Exposure of the LUX-ZEPLIN (LZ) Experiment J. Aalbers

### v1

@@ Abstract
Interactions in the LXe target that fills the TPC create light in the form of prompt vacuum ultraviolet (VUV) scintillation (S1) and delayed electroluminescence (S2), both observed by the top and bottom photomultiplier tube (PMT) arrays. The S2 signal is generated by ionization electrons as they traverse a thin layer of gaseous xenon above the liquid, having been drifted upwards from the interaction site under an applied electric field and extracted from the liquid phase. The measured time interval between the S1 and the S2 pulses, referred to as the “drift time”, corresponds to the depth, , of the interaction, while the hit pattern of S2 light in the top PMT array allows for position reconstruction in the transverse plane. The ratio of S2 to S1 signal sizes is used to discriminate between background electronic recoils (ERs) and nuclear recoils (NRs), the latter of which is expected for WIMP-nucleus scatters. Background events are further rejected if they exhibit multiple scatters (MS) rather than a single scatter (SS) in the TPC, or through the presence of coincident signals in the Skin and OD.
The LXe target is contained in a vacuum-insulated cryostat, and is continuously circulated for gas-phase purification to remove electronegative impurities, and to regulate the liquid level. The cooling and circulation systems allow for control of the LXe flow pattern inside the cryostat, which influences the spatial distribution and movement of radioisotopes in the TPC volume. Two distinct flow configurations can be attained, referred to as the “high-mixing” and “low-mixing” states. In the high-mixing state, the liquid flow results in spatially homogeneous distributions of injected calibration sources. In the low-mixing state, there exist distinct regions of laminar flow throughout the detector, and coincident 222RnPo decays can be reliably used to map the LXe velocity field. The ability to model the flow in this state is exploited in the development of a “radon tag” to target the prominent 214Pb ER background, discussed later.
@@
Two sources of background NR events are considered: those originating from neutrino interactions and those from radioactive decays in detector components. The neutrino interactions are from CENS of solar (8B and hep) and cosmic-ray-generated atmospheric neutrinos. Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and OD detectors. Events vetoed by these systems are used as a sideband that is used to derive an in-situ constraint on the number of neutron events in the WIMP search data. As discussed later, the sideband is incorporated directly into the final statistical analysis; however, a sideband-only fit returns an expectation of SS counts, as shown in Table 1. Secondary estimates from simulations based on detector material radioassays and from data using candidate MS neutron events tagged by the veto detectors return expectations of and neutron counts, respectively, consistent with the primary analysis.
@@
Two-neutrino DEC decays of 124Xe are observed in LZ with a half-life of [25], consistent with that in Ref. [26]. The expected number of 124Xe counts in the ROI is estimated assuming an abundance of [27] and branching ratios for the LL (10.00 keV) and LM (5.98 keV) decay modes of 1.4% and 0.8%, respectively. The LM charge suppression is modeled with the same ratio of 0.87 as used for single EC decays. The overlap of two L-shell cascades in LL-shell capture should result in additional recombination. As an independent measurement of the LL-shell charge yield is not available, the Thomas-Imel (TI) box model [28, 29] is invoked to relate the ionization density to the amount of electron-ion recombination at a given applied electric field. The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other by the expectation for a doubled ionization density from the TI box model (0.65).
@@
Figure 3: Black points show the final set of events passing all selection cuts with the panels distinguishing all events in WS2024 (top) from those that are radon tagged (bottom). 214Pb is at 60% and disperse backgrounds at 15% of their full FV activity in the radon-tagged data set; thus the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band. Dark and light, gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively. Green and orange contours show the same quantiles for the best-fit 124Xe distribution and the accidental coincidence events, respectively. Gray lines show contours of constant ER-equivalent (keVee) and NR-equivalent (keVnr) energy.
@@ Appendix A Appendix - Spin Dependent Results
Szydagis et al. [2022]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., arXiv preprint arXiv:2211.10726 (2022).
@@ Appendix C Details of the Detector Response Model
The LZ detector and xenon response models are implemented in a nest-based application that includes effects such as curved electron drift paths from field non-uniformities, finite position reconstruction resolution in the transverse and longitudinal directions, and position-dependence of the S1 and S2 signal sizes. The key detector parameters of the model are provided in Table S2.
@@
A header file for nest v2.4.0 that will reproduce the ER and NR response models used in this analysis is available online.

### v2

@@ Abstract
Interactions in the LXe target that fills the TPC create light in the form of prompt vacuum ultraviolet (VUV) scintillation (S1) and delayed electroluminescence (S2), both observed by the top and bottom photomultiplier tube (PMT) arrays. The S2 signal is generated by ionization electrons as they traverse a thin layer of gaseous xenon above the liquid, having been drifted upwards from the interaction site under an applied electric field and extracted from the liquid phase. The measured time interval between the S1 and the S2 pulses, referred to as the “drift time”, corresponds to the depth, , of the interaction, while the hit pattern of S2 light in the top PMT array allows for position reconstruction in the transverse plane. The ratio of S2 to S1 signal sizes is used to discriminate between background electronic recoils (ERs) and nuclear recoils (NRs), the latter of which is expected for WIMP-nucleus scatters. Background events are further rejected if they exhibit multiple scatters (MS) rather than a single scatter (SS) in the TPC, or through the presence of coincident signals in the Skin and/or OD.
The LXe target is contained in a vacuum-insulated cryostat, and is continuously circulated for gas-phase purification to remove electronegative impurities, and to regulate the liquid level. The cooling and circulation systems allow for control of the LXe flow pattern inside the cryostat, which influences the spatial distribution and movement of radioisotopes in the TPC volume. Two distinct flow configurations can be attained, referred to as the “high-mixing” and “low-mixing” states. In the high-mixing state, the liquid flow results in spatially homogeneous distributions of injected calibration sources. In the low-mixing state, there exist distinct regions of laminar flow throughout the detector, and coincident 222RnPo decays can be reliably used to map the LXe velocity field. The ability to model the flow in this state is exploited in the development of a “radon tag” to target the prominent 214Pb ER background (similar to that in Ref. [14]), discussed later.
@@
Two sources of background NR events are considered: those originating from neutrino interactions and those from radioactive decays in detector components. The neutrino interactions are from CENS of solar (8B and hep) and cosmic-ray-generated atmospheric neutrinos. Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and/or OD. Events vetoed by these systems are used as a sideband that is used to derive an in-situ constraint on the number of neutron events in the WIMP search data. As discussed later, the sideband is incorporated directly into the final statistical analysis; however, a sideband-only fit returns an expectation of SS counts, as shown in Table 1. Secondary estimates from simulations based on detector material radioassays and from data using candidate MS neutron events tagged by the veto detectors return expectations of and neutron counts, respectively, consistent with the primary analysis.
@@
Two-neutrino DEC decays of 124Xe are observed in LZ with a half-life of [26], consistent with that in Ref. [27]. The expected number of 124Xe counts in the ROI is estimated assuming an abundance of [28] and branching ratios for the LL (10.00 keV) and LM (5.98 keV) decay modes of 1.4% and 0.8%, respectively. The LM charge suppression is modeled with the same ratio of 0.87 as used for single EC decays. The overlap of two L-shell cascades in LL-shell capture should result in additional recombination. As an independent measurement of the LL-shell charge yield is not available, the Thomas-Imel (TI) box model [29, 30] is invoked to relate the ionization density to the amount of electron-ion recombination at a given applied electric field. The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other by the result from a preliminary study of increasing the ionization density in the TI box model (0.65).
@@
Figure 3: Black points show the final set of events passing all selection cuts with the panels distinguishing all events in WS2024 (top) from those that are radon tagged (bottom). The radon-tagged data set contains 60% and 15% of the full FV activity of the 214Pb and dispersed backgrounds, respectively; thus the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band. Dark and light, gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively. Green and orange contours show the same quantiles for the best-fit 124Xe distribution and the accidental coincidence events, respectively. Gray lines show contours of constant ER-equivalent (keVee) and NR-equivalent (keVnr) energy.
@@ Appendix A Appendix - Spin Dependent Results
Szydagis et al. [2022]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., arXiv:2211.10726 (2022).
@@ Appendix C Details of the Detector Response Model
[…]